OpenStudy (anonymous):

Solve. b/6+b+2/8≥-5/8 1.) b≤-9/7 2.) b≥9/7 3.) b≤ -3 4.) b≥ -3

OpenStudy (ajprincess):

Is this ur question? \[\frac{b}{6}+b+\frac{2}{8}\geq\frac{-5}{8}\]

OpenStudy (anonymous):

Yes

OpenStudy (ajprincess):

 Mutiply by 6 both sides and divide by 7 both sides. Can u do?